# Assessing Historical Structural Oppression Worldwide via Rule-Guided Prompting of Large Language Models

### v1

@@ Abstract
Traditional efforts to measure historical structural oppression struggle with cross-national validity due to the unique, locally specified histories of exclusion, colonization, and social status in each country, and often have relied on structured indices that privilege material resources while overlooking lived, identity-based exclusion. We introduce a novel framework for oppression measurement that leverages Large Language Models (LLMs) to generate context-sensitive scores of lived historical disadvantage across diverse geopolitical settings. Using unstructured self-identified ethnicity utterances from a multilingual COVID-19 global study, we design rule-guided prompting strategies that encourage models to produce interpretable, theoretically grounded estimations of oppression. We systematically evaluate these strategies across multiple state-of-the-art LLMs. Our results demonstrate that LLMs, when guided by explicit rules, can capture nuanced forms of identity-based historical oppression within nations. This approach provides a complementary measurement tool that highlights dimensions of systemic exclusion, offering a scalable, cross-cultural lens for understanding how oppression manifests in data-driven research and public health contexts. To support reproducible evaluation, we release an open-sourced benchmark dataset for assessing LLMs on oppression measurement (https://github.com/chattergpt/llm-oppression-benchmark).
@@ I Introduction
The study of racial and ethnic inequality remains central to sociological research, with extensive research documenting how structural oppression is reproduced in historical and contemporary contexts [1, 2, 3]. Oppression can be understood as a social hierarchy in which groups are relegated to lower status and subjected to systemic exclusion, dehumanization, and disadvantage. In public health and sociology, this is closely aligned with definitions of systemic and structural racism, which describe racism as deeply embedded in laws, policies, institutional practices, and social norms that sustain widespread inequities and oppression over time [1]. Foundational works have demonstrated how ethnic and national hierarchies shape access to power, life opportunities, autonomy, and sovereignty, for example, primarily through institutionalized mechanisms such as legal structures, educational systems, and healthcare access, among others [2]. These dynamics vary substantially across geopolitical regions, however, shaped by localized histories of colonization, racial and ethnic discrimination, and national classification systems  [4, 5]. As such, understanding racial and ethnic oppression defies cross-national analyses due to their somewhat idiosyncratic and varying origins, requiring methodologies that are sensitive to national contexts and histories. Traditional frameworks often rely on demographic indicators such as income, housing, and education, and perhaps standardized categories of race and/or ethnicity specific to that country, but can hardly capture the full spectrum of structural marginalization experienced by ethnic and racial groups across countries [6]. Additionally, oppression is a multidimensional phenomenon shaped not only by material violence but also by systemic exclusion from institutions, denial of access, fragility of social networks, and historically maintained systems of classification. Crucially, these processes are geographically and culturally contingent, varying across national boundaries and historical contexts [6, 7, 8]. Achieving health equity requires measures that go beyond material deprivation and account for dimensions such as structural racism, historical injustice, and cultural exclusion, which can be underrepresented in current indices [1]. Capturing these dimensions is especially challenging due to the localized nature of structural oppression, as each country has its own distinct history of exclusion, marginalization, and racial classification systems. A group that is marginalized in one society may be dominant in another, making it impossible to apply a single, global framework without oversimplifying these dynamics.
@@
Efforts to measure race and ethnicity in health and social research often rely on standardized categories imposed by government and statistical agencies, such as the Office of Management and Budget’s (OMB) minimum standards for race and ethnicity reporting in the United States (OMB15) [9, 10] or census classifications in other countries [11]. While these systems provide structured categories that enable comparability within national datasets, they frequently fail to capture how individuals actually identify [3]. Respondents are forced to choose categories that do not reflect their lived experience, and large umbrella groups such as “Asian” and “Hispanic” obscure substantial heterogeneity in histories, cultures, and health outcomes across subgroups. For example, aggregating South Asian, East Asian and Southeast Asian populations under a single “Asian” label masks wide disparities in disease risk, socioeconomic position and discrimination patterns [12]. Similarly, treating “Hispanic/Latino” as a homogeneous category overlooks differences across Mexican, Puerto Rican, Cuban and Central/South American populations in health outcomes and access to care [13]. These limitations highlight the core conflict between comparability and accuracy, as structured taxonomies facilitate statistical analysis but flatten identity-based variation, whereas free-text self-identification allows people to describe themselves in context-specific and intersectional ways but resists conventional quantitative analysis. These challenges underscore the need for new methodological approaches that can handle the flexibility of self-identified, unstructured identity data, motivating exploration of Large Language Models (LLMs) as tools for interpreting free-text inputs at scale.
@@
Large Language Models (LLMs) have emerged as tools for use in many domains, especially social sciences [14] and public health [15] to support tasks such as free-text extraction and classification, with high performance across a range of tasks when applied at scale [16]. However, among more sociological tasks, LLMs have presented their own issues and biases [17], including reproduction of racial and ethnic stereotypes and underrepresentation of structurally marginalized groups [18].
@@ II-A Quantifying Systemic Oppression Across Contexts
Currently, most efforts to quantify social disadvantage, such as the Index of Multiple Deprivation (IMD) in the UK [19] or regional deprivation indices in India [7] and Brazil [20], have relied on composite measures built from structured datasets, including census records, administrative data, and geospatial indicators [8, 21, 7], rather than flexible, open-ended utterances from people themselves. These tools, while valuable for resource allocation and public health monitoring, often suffer from three limitations. First, they privilege domains such as income and employment, often treating these metrics as identity-neutral and overlooking how race, ethnicity and structural power relations shape access to resources, thereby underrepresenting identity and power-based oppression dynamics. Second, they are often limited to national or subnational contexts and struggle with generalizability across borders. Third, their dependence on available administrative data often excludes experiences of discrimination that are not explicitly measured or encoded in existing survey methods [8, 21, 7].
@@
In addition to these limitations, standardized racial and ethnic categories used in census and administrative datasets are themselves politically derived and often carry inherent bias. For example, Native American respondents in the U.S. are often denied the opportunity to identify by tribal affiliation, despite the fact that tribal membership is central to accessing resources, legal status, and cultural identity [22]. These classifiers serve administrative or political purposes rather than reflecting the lived experiences of those being categorized [11]. In many cases, the very populations that experience systemic oppression are either excluded from official classifications altogether or absorbed under categories designed for bureaucratic convenience. By contrast, unstructured self-identification enables individuals to articulate their identities in their own terms, thereby democratizing ethnic categorization and revealing nuanced intersectional positions that structured indices overlook. This flexibility is particularly important for capturing marginalized, diasporic populations or mixed-identity populations whose self-descriptions may not neatly map onto standardized census categories.
@@
Developing oppression measurements requires careful selection of indicators grounded in theory, followed by standardization, weighting, and validation across contexts [6]. However, there is no consensus on how to assign relevant domains in global or multi-ethnic contexts. Factor analysis techniques like Principal Component Analysis (PCA) and its geographically weighted variant (GWPCA) are often used to derive scores, but these approaches struggle with replicability over time. Additionally, while validation is commonly performed against health outcomes or existing indices, such efforts rarely disentangle the specific role of identity-based oppression from broader measures of material disadvantage [7, 8].
@@
These limitations reflect broader concerns in the biomedical and social science literature regarding the misuse and oversimplification of race and ethnicity in research, and the denial of autonomy in self-identifying one’s ethnicity. Race and ethnicity are frequently treated as static variables without sufficient consideration of their sociopolitical construction, contextual meanings, or structural forces that shape differential health outcomes. The National Academies of Sciences, Engineering, and Medicine have called for researchers to adopt principled, theory-driven definitions of race and ethnicity, disclose data provenance, and critically assess the rationale and limitations behind their use [23]. In alignment with these recommendations, our study conceptualizes oppression as an outcome of institutional power and historical injustice rather than simply a demographic descriptor.
@@ II-B LLMs and Sociological Reasoning
An emerging area of application for LLMs is survey research, where they show promise for generating synthetic responses, clarifying question-wording, and mitigating response bias by simulating varied interpretations of survey items [24, 25]. These applications highlight how LLMs can be integrated into existing workflows to improve data quality and enhance interpretability with methodological care and rigor.
However, few studies have evaluated whether LLMs can produce valid, replicable judgments on concepts such as discrimination, oppression, or privilege, especially when grounded in sociological theory. Recent evaluations have primarily focused on fairness in text generation, with limited attention to how LLMs approximate sociological indices [26, 27, 28, 17]. Additionally, existing tools to quantify systemic oppression struggle with scalability and cross-national generalization, often relying on expert-coded taxonomies or language-specific analyses [6, 7, 8], or inappropriate adoption of one country’s predominant categories in another, which can effectively export one country’s racist construction of ethnicity to another country. This limitation underscores the need for flexible approaches that are adaptable across cultural and linguistic concepts, motivating our use of instruction-tuned LLMs guided by sociologically-informed prompts and scoring rubrics.
@@ III-A Schema Construction
We adopt a bottom-up approach to construct a schema for identity-based historical oppression classification. Our starting point is a multilingual global survey on COVID-19 non-medical COVID-related impact [29], where participants were asked to answer the open-ended question “How would you describe your ethnic background?” (other language variants are provided in Appendix -A1) under the Demographic Questions section of the survey to self-identify their ethnic group. Participants were also asked to choose their country of residence from a list of over 170 countries, with an option for self-description if their country was not provided. This is a decision that prioritizes each person’s current social context – regardless of the meaning of their identity in their birth country (if different) – which can change dramatically in changing from one country of residence to another. The resulting dataset contains a wide range of free-text responses across multiple languages (e.g., English, Spanish, Italian, Mandarin, Arabic, ASL, French, Hindi), offering a diverse window into global identity categories. Unlike traditional census methods for identifying race or ethnicity, these free-text descriptions allow study participants to identify themselves and their own identities rather than choosing from pre-defined categories.
@@
We review and cluster these responses, guided by critical theoretical frameworks including cumulative disadvantage [30], racial formation [23], historical trauma [31], and critical race theory [2, 32]. From this analysis, we develop a five-level ordinal classification that conceptualizes oppression as a group’s institutional positioning, historical marginalization, and exposure to structural violence. The schema captures a comprehensive continuum and accommodates all observed responses:
@@ III-B1 Prompt Template and Components
Prompts are assembled in the following order: (1) system role $\rightarrow$ (2) identity statement $\rightarrow$ (3) instruction $\rightarrow$ (4) oppression schema $\rightarrow$ (5) required output format.
@@ III-B2 Rule-Guided Prompt Module
These rules provide a principled scaffold for LLM reasoning, ensuring alignment with historical and sociological frameworks. For instance, Rule 2 enables differentiation of “Hispanic” respondents in Brazil versus the United States, while Rule 4 acknowledges the impact of the historically pervasive “one-drop rule” to correctly classify mixed-ancestry groups. The “one-drop” was a legal and social doctrine in the United States that classified any individual with a trace of African ancestry as Black, and was codified in state statutes and upheld as a tool to enforce rigid racial boundaries during the Jim Crow Era [33]. We reference the rule only as a historical example of ancestry-based exclusion, not as an endorsement. While the one-drop rule is specific to the United States, similar logics of exclusion based on partial marginalized ancestry exist in other countries, albeit with different historical foundations. For example, in Brazil and other Latin American countries, individuals with mixed African ancestry may be socially classified in less marginalized categories depending on skin color or phenotype, yet still encounter discrimination tied to their ancestry [34]. Additionally, Rule 3 ensures consistency when national identifiers are used, defaulting to dominant groups unless modifiers indicate otherwise.
@@ IV-A1 Baselines
To demonstrate the contribution of the rule-guided prompt module, we compare against two prompting strategies:
@@ IV-A2 Language Models
  • Gemini 1.5 Pro [35]: A multimodal model developed by Google DeepMind.
  • GPT-3.5 Turbo [36]: An instruction-tuned model created by OpenAI.
  • GPT-4o mini [37]: A fast, low-latency variant of GPT-4 released by OpenAI.

### v2

@@ Abstract
Traditional efforts to measure historical structural oppression struggle with cross-national validity due to the unique, locally specified histories of exclusion, colonization, and social status in each country, and often have relied on structured indices that privilege material resources while overlooking lived, identity-based exclusion. We introduce a novel framework for oppression measurement that leverages Large Language Models (LLMs) to generate context-sensitive scores of lived historical disadvantage across diverse geopolitical settings. Using unstructured self-identified ethnicity utterances from a multilingual COVID-19 global study, we design rule-guided prompting strategies that encourage models to produce interpretable, theoretically grounded estimations of oppression. We systematically evaluate these strategies across multiple state-of-the-art LLMs. Our results demonstrate that LLMs, when guided by explicit rules, can capture nuanced forms of identity-based historical oppression within nations. This approach provides a complementary measurement tool that highlights dimensions of systemic exclusion, offering a scalable, cross-cultural lens for understanding how oppression manifests in data-driven research and public health contexts. To support reproducible evaluation, we release an open-sourced benchmark dataset for assessing LLMs on oppression measurement (https://github.com/chattergpt/HSO-Bench).
@@ I Introduction
The study of racial and ethnic inequality remains central to sociological research, with extensive research documenting how structural oppression is reproduced in historical and contemporary contexts [braveman2021, bonillasilva1997, Saperstein2013]. Oppression can be understood as a social hierarchy in which some groups subject other groups to lower status and to systemic exclusion, dehumanization, and disadvantage. In public health and sociology, this oppression is closely aligned with definitions of systemic and structural racism, which describe racism as deeply embedded in laws, policies, institutional practices, and social norms that sustain widespread inequities, violence, and disadvantage over time [braveman2021]. Foundational works have demonstrated how ethnic and national hierarchies shape access to power, life opportunities, autonomy, and sovereignty, for example, primarily through institutionalized mechanisms such as legal structures, educational systems, and healthcare access, among others [bonillasilva1997]. These dynamics vary substantially across geopolitical regions, however, shaped by localized histories of colonization, racial and ethnic discrimination, and national classification systems  [brubaker2006, hall2000]. As such, understanding racial and ethnic oppression defies cross-national analyses due to their somewhat idiosyncratic and varying origins, requiring methodologies that are sensitive to national contexts and histories. Traditional frameworks often rely on demographic indicators such as income, housing, and education, and perhaps standardized categories of race and/or ethnicity specific to that country, but can hardly capture the full spectrum of structural marginalization experienced by ethnic and racial groups across countries [allik2020]. Additionally, oppression is a multidimensional phenomenon shaped not only by material violence but also by systemic exclusion from institutions, denial of access, fragility of social networks, and historically maintained systems of classification. Crucially, these processes are geographically and culturally contingent, varying across national boundaries and historical contexts [allik2020, basu2020, wang2021]. Achieving health equity requires measures that go beyond material deprivation and account for dimensions such as structural racism, historical injustice, and cultural exclusion, which can be underrepresented in current indices [braveman2021]. Capturing these dimensions is especially challenging due to the localized nature of structural oppression, as each country has its own distinct history of exclusion, marginalization, and racial classification systems. A group that is marginalized in one society may be dominant in another, making it impossible to apply a single, global framework without oversimplifying these dynamics.
@@
Efforts to measure race and ethnicity in health and social research often rely on standardized categories imposed by government and statistical agencies, such as the Office of Management and Budget’s (OMB) minimum standards for race and ethnicity reporting in the United States (OMB15) [OMB1997, OMB2016] or census classifications in other countries [Morning2008]. While these systems provide structured categories that enable comparability within national datasets, they frequently fail to capture how individuals actually identify [Saperstein2013]. Respondents are forced to choose categories that do not reflect their lived experience, and large umbrella groups such as “Asian” and “Hispanic” obscure substantial heterogeneity in histories, cultures, and health outcomes across subgroups. For example, aggregating South Asian, East Asian, and Southeast Asian populations under a single “Asian” label masks wide disparities in disease risk, socioeconomic position and discrimination patterns [Islam2010]. Similarly, treating “Hispanic/Latino” as a homogeneous category overlooks differences across Mexican, Puerto Rican, Cuban, and Central/South American populations in health outcomes and access to care [Lara2005]. These limitations highlight the core conflict between comparability and accuracy, as structured taxonomies facilitate statistical analysis but flatten identity-based variation, whereas free-text self-identification allows people to describe themselves in context-specific and intersectional ways but resists conventional quantitative analysis. These challenges underscore the need for new methodological approaches that can handle the flexibility of self-identified, unstructured identity data, motivating exploration of Large Language Models (LLMs) as tools for interpreting free-text inputs at scale.
@@
Large Language Models (LLMs) have emerged as tools for use in many domains, especially social sciences [lyu2025gpt, mou2024unifying] and public health [jo2023understanding] to support tasks such as free-text extraction and classification, with high performance across a range of tasks when applied at scale [guo2024llmhealth]. However, among more sociological tasks, LLMs have presented their own issues and biases [qi2025representation], including reproduction of racial and ethnic stereotypes and underrepresentation of structurally marginalized groups [sheng2021bias].
@@ II-A Quantifying Systemic Oppression Across Contexts
Currently, most efforts to quantify social disadvantage, such as the Index of Multiple Deprivation (IMD) in the UK [imd2019] or regional deprivation indices in India [basu2020] and Brazil [allik2020brazdep], have relied on composite measures built from structured datasets, including census records, administrative data, and geospatial indicators [wang2021, maroko2016, basu2020], rather than flexible, open-ended utterances from people themselves. These tools, while valuable for resource allocation and public health monitoring, often suffer from three limitations. First, they privilege domains such as income and employment, often treating these metrics as identity-neutral and overlooking how race, ethnicity and structural power relations shape access to resources, thereby underrepresenting identity and power-based oppression dynamics. Second, they are often limited to national or subnational contexts and struggle with generalizability across borders. Third, their dependence on available administrative data often excludes experiences of discrimination that are not explicitly measured or encoded in existing survey methods [wang2021, maroko2016, basu2020].
@@
In addition to these limitations, standardized racial and ethnic categories used in census and administrative datasets are themselves politically derived and often carry inherent bias. For example, Native American respondents in the U.S. are often denied the opportunity to identify by tribal affiliation, despite the fact that tribal membership is central to accessing resources, legal status, and cultural identity [Snipp2003]. These classifiers serve administrative or political purposes rather than reflecting the lived experiences of those being categorized [Morning2008]. In many cases, the very populations that experience systemic oppression are either excluded from official classifications altogether or absorbed under categories designed for bureaucratic convenience. By contrast, unstructured self-identification enables individuals to articulate their identities in their own terms, thereby democratizing ethnic categorization and revealing nuanced intersectional positions that structured indices overlook. This flexibility is particularly important for capturing marginalized, diasporic populations or mixed-identity populations whose self-descriptions may not neatly map onto standardized census categories.
@@
Developing oppression measurements requires careful selection of indicators grounded in theory, followed by standardization, weighting, and validation across contexts [allik2020]. However, there is no consensus on how to assign relevant domains in global or multi-ethnic contexts. Factor analysis techniques like Principal Component Analysis (PCA) and its geographically weighted variant (GWPCA) are often used to derive scores, but these approaches struggle with replicability over time. Additionally, while validation is commonly performed against health outcomes or existing indices, such efforts rarely disentangle the specific role of identity-based oppression from broader measures of material disadvantage [basu2020, wang2021].
@@
These limitations reflect broader concerns in the biomedical and social science literature regarding the misuse and oversimplification of race and ethnicity in research, and the denial of autonomy in self-identifying one’s ethnicity. Race and ethnicity are frequently treated as static variables without sufficient consideration of their sociopolitical construction, contextual meanings, or structural forces that shape differential health outcomes. The National Academies of Sciences, Engineering, and Medicine have called for researchers to adopt principled, theory-driven definitions of race and ethnicity, disclose data provenance, and critically assess the rationale and limitations behind their use [powe2025race]. In alignment with these recommendations, our study conceptualizes oppression as an outcome of institutional power and historical injustice rather than simply a demographic descriptor.
@@ II-B LLMs and Sociological Reasoning
An emerging area of application for LLMs is survey research, where they show promise for generating synthetic responses, clarifying question-wording, and mitigating response bias by simulating varied interpretations of survey items [jansen2023survey, zhang2025socioverse]. These applications highlight how LLMs can be integrated into existing workflows to improve data quality and enhance interpretability with methodological care and rigor.
However, few studies have evaluated whether LLMs can produce valid, replicable judgments on concepts such as discrimination, oppression, or privilege, especially when grounded in sociological theory. Recent evaluations have primarily focused on fairness in text generation, with limited attention to how LLMs approximate sociological indices [gebru2021datasheets, srivastava2022beyond, liang2022holistic, qi2025representation]. Additionally, existing tools to quantify systemic oppression struggle with scalability and cross-national generalization, often relying on expert-coded taxonomies or language-specific analyses [allik2020, basu2020, wang2021], or inappropriate adoption of one country’s predominant categories in another, which can effectively export one country’s racist construction of ethnicity to another country. This limitation underscores the need for flexible approaches that are adaptable across cultural and linguistic concepts, motivating our use of instruction-tuned LLMs guided by sociologically-informed prompts and scoring rubrics.
@@ III-A Schema Construction
We adopt a bottom-up approach to construct a schema for identity-based historical oppression classification. We adopt a bottom-up approach to construct a schema for identity-based historical oppression classification. Our starting point is a secondary analysis of a multilingual global survey on COVID-19 non-medical COVID-related impact [dye2021covid], where participants were asked to answer the open-ended question “How would you describe your ethnic background?” (other language variants are provided in Appendix -A1) under the Demographic Questions section of the survey to self-identify their ethnic group. Participants were also asked to choose their country of residence from the standardized list of countries provided by International Standard Organization 3166 Country Codes, with an option for self-description if their country was not provided [iso_obp]. This is a decision that prioritizes each person’s current social context – regardless of the meaning of their identity in their birth country (if different) – which can change dramatically in changing from one country of residence to another. The resulting dataset contains a wide range of free-text responses across multiple languages (e.g., English, Spanish, Italian, Mandarin, Arabic, ASL, French, Hindi), offering a diverse window into global identity categories. Unlike traditional census methods for identifying race or ethnicity, these free-text descriptions allow study participants to identify themselves and their own identities rather than choosing from pre-defined categories.
@@
We review and cluster these responses, guided by critical theoretical frameworks including cumulative disadvantage [dannefer2003cumulative], racial formation [powe2025race], historical trauma [braveheart1998american], and critical race theory [bonillasilva1997, pierson2004politics]. From this analysis, we develop a five-level ordinal classification that conceptualizes oppression as a group’s institutional positioning, historical marginalization, and exposure to structural violence. The scale is theory-driven, as we do not assume equal intervals between categories and is not empirically calibrated via qualitative outcome data. Rather, it is derived from historical and sociopolitical context, reflecting increasing degrees of systemic exclusion and marginalization. The schema captures a comprehensive continuum and accommodates all observed responses:
@@ III-B1 Prompt Template and Components
Prompts are assembled in the following order: (1) system role $\rightarrow$ (2) identity statement $\rightarrow$ (3) instruction $\rightarrow$ (4) oppression schema $\rightarrow$ (5) required output format. This is illustrated in Figure 1.
@@ III-B2 Rule-Guided Prompt Module
These rules provide a principled scaffold for LLM reasoning, ensuring alignment with historical and sociological frameworks. For instance, Rule 2 enables differentiation of “Hispanic” respondents in Brazil versus the United States, while Rule 4 acknowledges the impact of the historically pervasive “one-drop rule” to correctly classify mixed-ancestry groups. The “one-drop” was a legal and social doctrine in the United States that classified any individual with a trace of African ancestry as Black, and was codified in state statutes and upheld as a tool to enforce rigid racial boundaries during the Jim Crow Era [Khanna2010]. We reference the rule only as a historical example of ancestry-based exclusion, not as an endorsement. While the one-drop rule is specific to the United States, similar logics of exclusion based on partial marginalized ancestry exist in other countries, albeit with different historical foundations. For example, in Brazil and other Latin American countries, individuals with mixed African ancestry may be socially classified in less marginalized categories depending on skin color or phenotype, yet still encounter discrimination tied to their ancestry [Telles2014]. Additionally, Rule 3 ensures consistency when national identifiers are used, defaulting to dominant groups unless modifiers indicate otherwise.
@@ IV Experiments
Figure 1: Illustration of the prompting setup used for our experiment, comparing Rule-Guided Prompting (RG) with two baseline methods: Vanilla and Chain-of-Thought (CoT). Each prompt is constructed sequentially as follows: (1) system role $\rightarrow$ (2) identity statement $\rightarrow$ (3) instruction $\rightarrow$ (4) oppression schema $\rightarrow$ (5) optional CoT or RG instruction $\rightarrow$ (6) required output format. The complete set of rules for the RG method is provided in Section III-B2, and the set of guiding questions for the CoT method is detailed in Section IV-A1.
@@ IV-A1 Baselines
To demonstrate the contribution of the rule-guided prompt module, we compare against two prompting strategies, which is illustrated in Figure 1:
@@ IV-A2 Language Models
  • Gemini 1.5 Pro [team2024gemini]: A multimodal model developed by Google DeepMind.
  • GPT-3.5 Turbo [ouyang2022training]: An instruction-tuned model created by OpenAI.
  • GPT-4o mini [hurst2024gpt]: A fast, low-latency variant of GPT-4 released by OpenAI.